A physics teacher sits on a rotating stool to demonstrate the conversion of angular motion. with her arms stretched, she rotates at 0.75 rev/s. When teacher's arms are tucked in close in her chest, she rotates at 1.50 rev/s. What is the ratio of her moment of inertia with outstretched arms to the moment of inertia with arms tucked in to her chest?
1
Expert's answer
2020-12-15T11:47:07-0500

Let's write the law of conservation of angular momentum:


"L_{i}=L_{f},""I_{str.}\omega_{str.}=I_{tuck.}\omega_{tuck.},""\dfrac{I_{str.}}{I_{tuck.}}=\dfrac{\omega_{tuck.}}{\omega_{str.}}=\dfrac{1.50\ \dfrac{rev}{s}}{0.75\ \dfrac{rev}{s}}=2."

Answer:

"\dfrac{I_{str.}}{I_{tuck.}}=2."